OpenStudy (anonymous):

Solve the proportion: 16/17=x/40 Solve the proportion: 16/17=x/40 @Mathematics

OpenStudy (anonymous):

A. 42.5 B. 37.6 C. -42.5 D. -37.6

jimthompson5910 (jim_thompson5910):

16/17 = x/40 16*40 = 17x 640 = 17x 640/17 = x 37.6470588235294 = x x = 37.6470588235294 Now round to the nearest tenth to get x = 37.6 So the answer is choice B